# MicroDreamer: Zero-shot 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction

### v1

@@ 4 MicroDreamer
5 Experiment
We present the results on NeRF, 3DGS, and ablation sequentially. For a fair comparison, we utilize the widely used CLIP similarity [41] for quantitative comparison following DreamGaussian [55]. We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0. For all methods, the corresponding images are rendered from NeRF in Sec. 5.1 and mesh in Sec. 5.2. We report the average CLIP similarities between these images and the reference image or text. See Appendix C for details of the checkpoint and codebase.
5.1 Results on NeRF
We apply SIR algorithm on NeRF[34, 36, 9] leveraging three multi-view diffusion model from MVDream[47], Stable Zero123[52, 24], and ImageDream[58]. For each model, we selected 6 input prompts from a widely used codebase [9] for testing, calculated NFEs, and recorded the change of the average CLIP similarity in the generation process. Compared with SDS, SIR lowers the total NFEs while generating a 3D object and can accelerate the generation process 5-20 times, as shown in Fig. 3. Besides, SIR can generate NeRF of high visual quality, detailed in Appendix D, and the conclusion agrees with the results in Fig. 3.
(a) Comparsion w.r.t. NFEs.
(b) Comparison on MVDream[47]
(c) Comparison on Stable Zero123[52, 24]
(d) Comparison on ImageDream[58]
Figure 3: Comparison of SIR and SDS on NeRF. (a) shows that SIR lowers the total NFEs. (b-d) show the change of the CLIP similarity in the generation process under different models and SIR achieves a 5-20 times acceleration approximately.
5.2 Results on 3D Gaussian Splatting
Qualitative comparisons.
We show the generated 3D objects amongst various methods and ours in Fig. 4. The baselines include representative feed-forward models One-2-3-45 [23] and TriplaneGaussian [67], and the fastest optimization-based model DreamGaussian [55]. We present the exported mesh from these models. MicroDreamer generates higher quality mesh exported from 3DGS than the feed-forward baselines and halves the time of DreamGaussian [55] while retaining competitive performance.
Figure 4: Qualitative comparisons on mesh from 3DGS. MicroDreamer achieves an excellent efficiency-performance balance in terms of the visual quality of the mesh.
Quantitative comparisons.
We compare MicroDreamer with seven competitive baselines including Point-E [37], Shap-E [14], One-2-3-45 [23], TriplanceGaussian [67], Wonder3D [27], LGM [54] and DreamGaussian [55]. We record the optimization time  [^3] The detailed time recording method is provided in Appendix C. for each model on a single NVIDIA A100 (80GB) GPU, as shown in Fig. 5(a), and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works[23, 25, 55], as shown in Fig. 5(b). As a zero-shot method, MicroDreamer is on par with speed compared to feed-forward methods, faster than DreamGaussian, with a very competitive CLIP similarity. We provide detailed tables in Appendix E.
Notably, the sample quality of MicroDreamer can be improved as the multi-view diffusion evolves. See detailed limitations and future work in Sec. 7.
(a) Optimization time for mesh on a single A100.
(b) CLIP similarity versus inference time.
Figure 5: Quantitative comparisons: efficiency versus quality. MicroDreamer achieves an excellent efficiency-effectiveness balance in terms of the CLIP similarity [41]. The optimization-based methods are highlighted in red in (b).
5.3 Ablation Study
(a) Optimization space [52, 24].
(b) Optimization space [58].
(c) Iterative reconstruction.
(d) Number of camera views.
(e) Time schedule.
(f) Forward process.
Figure 6: Ablation study. (a-b) Optimization in pixel space accelerates the generation process. (c) Iterative reconstruction is necessary. (d) More camera views benefit. (e) A linear schedule is sufficient. (f) A hybrid forward process can be effective.
Ablation on optimization space.
From the results in Fig. 6(a) and Fig. 6(b), we can see that SIR-pixel achieves 2-3 times more acceleration than SIR-latent, verifying the benefits of SIR on enabling optimization in the pixel space in Sec. 3.3.
Ablation on iterative reconstruction.
We highlight the necessity of iterative reconstruction in Fig. 6(c). The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise. Although increasing $K$ improves the quality, we set $K\leq 30$ for efficiency.
Ablation on number of camera poses.
SIR necessitates more than one camera pose for reconstruction, as shown in Fig. 6(d). The performance increases with more camera poses. We choose $6$ on 3DGS with the Stable Zero123 [52] model to balance efficiency and quality. See Tab. 1 for the values in other settings.
Ablation on time schedule.
We compare our linear schedule, the random schedule employed in SDS [40], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.8T-0.2T)+0.2T$. As shown in Fig. 6(e), the schedules perform similarly. For simplicity and efficiency, we use the linear schedule by default.
Ablation on forward process.
Fig. 6(f) compares three forward processes on 3DGS with the Stable Zero123 [52] model. The hybrid strategy performs better than noise-adding and is more efficient than inversion as discussed in Sec. 3.2. See Tab. 1 for other settings.
6 Related work
Optimization-based 3D Generation.
Built upon text-to-image diffusion models, optimization-based approaches [40, 56, 19, 66, 5, 24, 46, 60, 55, 18, 53, 33, 62, 23, 22] generate 3D objects without additional 3D data. Among them, the most relevant line of work [40, 60, 18, 66, 56] proposes various distillation algorithms. Besides SDS [40], ProlificDreamer [60] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference. LucidDreamer [18] incorporates DDIM inversion to strengthen the forward process. As mentioned in Sec. 1, though promising, such methods suffer from inefficiency problems due to the large number of function evaluations (NFE) and optimization in the latent space. Distinct from existing algorithms, SIR mimics the 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion and enables optimization in the pixel space. Our experiments demonstrate that SIR is 5-20 times faster than SDS [40] on NeRF and twice faster than the most competitive baseline DreamGaussian [55] on meshes.
Feed-forward 3D Generation.
Contrary to optimization-based methods, feed-forward approaches[3, 4, 6, 14, 37, 26, 13, 35, 59, 61, 65, 67] elect to use large-scale 3D datasets[8, 7] training to achieve a process that can directly generate 3D objects. These methods are characterized by their speed but they often produce lower-quality 3D structures with simple textures. Several recent methods, utilizing more data, have exhibited possibilities. For instance, the large reconstruction model (LRM)[10, 13] demonstrates the feasibility of training a Transformer to achieve a NeRF from a single image. Instant3D[17] further expands on the LRM enabling it to accept multi-view input and reconstruct it into a NeRF. The recent work large multi-view Gaussian model (LGM)[54] employs a similar concept, generating 3DGS from multi-view images. Remarkably, our MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [41, 55].
Multi-view Prediction based 3D Generation.
There is also a line of work[27, 25] dedicated to enhancing the output of multi-view diffusion models to reconstruct 3D objects. Such systems, like SyncDreamer[25] and Wonder3D[27], train or fine-tune multi-view diffusion models on 3D datasets and use a single reconstruction process to obtain 3D objects from images without iteration. These methods typically rely on 3D prior knowledge and are limited to specific 3D representations [25, 27, 57] with a long reconstruction time. In contrast, thanks to the carefully designed iterative process in SIR, MicroDreamer applies to various 3D representations and is much more efficient (see Fig. 5(a) for a direct comparison).
7 Conclusion
We introduce SIR, an efficient and general algorithm mimicking the 3D reconstruction process to reduce the total NFEs in zero-shot 3D generation. Based on SIR, our method MicroDreamer achieves a 5 to 20 times speed increase for NeRF generation compared to SDS on various base models. Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest zero-shot methods and matching the speed of feed-forward approaches, with a competitive generation quality.
Limitations.
There is potential for further improving the efficiency of MicroDreamer via employing consistency models [50, 32] or alternative sampling models that require fewer steps [20, 63, 45]. Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion. Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
Potential Negative Impact.
As a generative approach, the paper can potentially be used for fabricating data and news. Moreover, the 3D content generated from input images may infringe on the privacy or copyright of others. While automatic detection might mitigate these issues, they warrant careful consideration.
8 Acknowledgement
Appendix 0.A Results of Applying SDS in Pixel Space.
@@ 8 Acknowledgement
We highlight that SDS optimizing in pixel space can not be effective even if it has the data predicting form by reparameterized as presented in Eq. 6. Fig. 7 shows the poor quality for single-step prediction from the diffusion model using Eq. 2. Therefore, if we attempt to employ SDS for optimization in the pixel space, i.e. optimize 3D objects with 2D images like those generated in Fig. 7, the resulting 3D content would be of subpar quality, as the results presented in Fig. 8.
@@
Appendix 0.B Time Schedule of $t_{1}$
The hyperparameter $t_{1}$ regulates the balance between inversion and noise-adding, which varies with different models. In our earlier exploration, we adopted $t_{1}=t^{2}_{2}/T$ for SIR based on MVDream [47] and ImageDream [58] on NeRF, and $t_{1}=0.6t_{2}$ for SIR based on Stable Zero123 [52, 24] on both NeRF and 3DGS. Notably, SIR based on ImageDream on 3DGS opted for $t_{1}=t_{2}$, i.e., adding noise for the forward process. As an advanced multi-view diffusion model, ImageDream can generate images with enhanced 3D consistency, obviating the need for inversion to maintain consistency. Fig. 9 shows some results of SIR for ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
@@ Appendix 0.B Time Schedule of $t_{1}$
Appendix 0.C More Implementation Details
We present the codebase, some details in metrics, and the data used in ablation sequentially.
0.C.1 Codebase
For the implementation of SIR and SDS on NeRF, we choose the popularly used framework threestudio [9]. For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [55] [^4] https://github.com/dreamgaussian/dreamgaussian. The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [^5] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k.
In Sec. 5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub, including Point-E [37] [^6] https://github.com/openai/point-e/, Shap-E [14] [^7] https://github.com/openai/shap-e, One-2-3-45 [23] [^8] https://github.com/One-2-3-45/One-2-3-45, TriplanceGaussian [67] [^9] https://github.com/VAST-AI-Research/TriplaneGaussian, Wonder3D [27] [^10] https://github.com/xxlong0/Wonder3D, LGM [54] [^11] https://github.com/3DTopia/LGM and DreamGaussian [55] [^12] https://github.com/dreamgaussian/dreamgaussian.
0.C.2 Details in Metric
All the results presented in Sec. 5.2 were on the mesh representation. As TriplaneGaussian [67] has no official mesh export code, we apply the mesh exported code from LGM [54] for it.
For extra refine time, we record the time of second stage refinement for MicroDreamer and DreamGaussian [55], the time of generating 3D from multi-view images for Wonder3D[27], and the time for meshing through a NeRF for LGM[54], respectively.
0.C.3 Data Used in Ablation Study
We chose seven images collected from the test dataset to perform our ablation study. We will include these images in the submitted supplementary material.
Appendix 0.D More Results on NeRF
@@
Appendix 0.E Quantative Comparison on 3DGS
We provide the detailed quantitative comparison table in Tab. 2. For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iter. As shown in Tab. 2, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
@@ Appendix 0.E Quantative Comparison on 3DGS
DreamGaussian-300 iter 0.656 $\sim 18$s $\sim 10$s
MicroDreamer-20 iter (Ours) 0.661 $\sim 18$s $\sim 4$s
MicroDreamer-30 iter (Ours) 0.687 $\sim 26$s $\sim 4$s

### v2

@@ Appendix A Key hyperparameters of MicroDreamer
Appendix B The impact of latent diffusion model on 3D generation
Instead of applying the diffusion model directly in the pixel space to generate images, existing research [17] tends to use a VAE [75, 76] encoder to compress high-dimensional image data into a low-dimensional latent space first. Then a diffusion process is employed within this latent space to generate new images, which are finally decoded back by a VAE decoder to the pixel space to achieve high-quality images. We refer to this type of diffusion model as the latent diffusion model (LDM). The framework of LDM is adopted widely [6, 42, 44, 43], which leads to a significant impact on 3D generation.
B.1 SIR enabling optimization in the pixel space
Within the framework of LDM [17], the rendered images are mapped through an encoder $\mathcal{E}$ [75, 76] to a latent space, where the loss function is calculated. Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation. In such a case, the SIR loss in the latent space is given by:
However, we argue that this is not feasible for existing optimization methods [1, 2, 8] such as SDS. Even reparameterized as the data predicting form in Eq. (6), the pixel-level quality of the images obtained by SDS is poor for 3D optimization (see Appendix B.2). In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space. The SIR loss in the pixel space is formalized as:
which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as the ablation results shown in Fig. 8. It is set as the default loss throughout the paper.
Figure 8: Optimization in pixel space accelerates generation.
B.2 Faliure of applying SDS in pixel space.
@@ B.1 SIR enabling optimization in the pixel space
We highlight that SDS optimizing in pixel space can not be effective even if it has the data predicting form by reparameterized as presented in Eq. (6). Fig. 9 shows the poor quality for single-step prediction from the diffusion model using Eq. (2). Therefore, if we attempt to employ SDS for optimization in the pixel space, i.e. optimizing 3D objects with 2D images like those generated in Fig. 9, the resulting 3D content would be of subpar quality, as the results presented in Fig. 10.
@@
Appendix C Time schedule of $t_{1}$
The hyperparameter $t_{1}$ regulates the balance between inversion and noise-adding, which varies with different models. In our earlier exploration, we adopted $t_{1}=t^{2}_{2}/T$ for SIR based on MVDream [42] and ImageDream [44] on NeRF, and $t_{1}=0.6t_{2}$ for SIR based on Stable Zero123 [43, 6] on both NeRF and 3DGS. Notably, SIR based on ImageDream on 3DGS opted for $t_{1}=t_{2}$, i.e., adding noise for the forward process. As an advanced multi-view diffusion model, ImageDream can generate images with enhanced 3D consistency, obviating the need for inversion to maintain consistency. Fig. 11 shows some results of SIR on ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
@@ Appendix C Time schedule of $t_{1}$
Appendix D More implementation details
Codebase.
For the implementation of SIR and SDS on NeRF, we choose the popularly used framework threestudio [56]. For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [9]. The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [45, 77, 78]. In Sec. 5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub. See Tab. 2 for details about the URL and licenses of the assets we used in this paper.
Table 2: Codebase. URL and licenses are provided for the open-source assets we used in this paper.
Details in metric.
All the results presented in Sec. 5.2 were on the mesh representation. As TriplaneGaussian [33] has no official mesh export code, we apply the mesh exported code from LGM [30] for it. For extra refine time, we record the time of second stage refinement for MicroDreamer and DreamGaussian [9], the time of generating 3D from multi-view images for Wonder3D [39], and the time for meshing through a NeRF for LGM [30], respectively.
Data used in ablation study.
We chose seven images collected from the test dataset to perform our ablation study in Fig. 7 and Fig. 8. We will include these images in the submitted supplementary material.
Safeguards.
Our method, which requires no additional models and doesn’t release any new pre-trained model, can ensure security due to the inherent safe configuration of the employed 2D multi-view diffusion models [42, 43, 44].
Appendix E More results
Results on 3DGS.
We provide more results on mesh exported from 3DGS for MicroDreamer-20 iters in Fig. 12.
Figure 12: More results for MicroDreamer-20 iters.
Results on NeRF.
@@ Appendix E More results
Limitations.
In Fig. 14, we show some of the less successful cases generated by our method. Due to the limited ability of the diffusion model, inputs with complex geometric structures such as central hollowing may not result in satisfactory mesh outcomes. Furthermore, existing diffusion models struggle to generate the rear view of an object, which often leads to poorly textured back surfaces of the generated meshes. These problems may be solved as the quality of the images the multi-view diffusion model generated improves.
Figure 14: Limitations. Visualization of some less satisfactory cases generated by our methods.
Appendix F Quantative comparison on 3DGS
We provide the detailed quantitative comparison table in Tab. 3. For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iters. As shown in Tab. 3, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
@@ Appendix F Quantative comparison on 3DGS
DreamGaussian-300 iters 0.656 $\sim 18$s $\sim 10$s
MicroDreamer-20 iters (Ours) 0.678 $\sim 18$s $\sim 4$s
MicroDreamer-30 iters (Ours) 0.687 $\sim 26$s $\sim 4$s